# RankAlign: A Ranking View of the Generator-Validator Gap in Large Language Models

### v1

@@ Abstract
Although large language models (LLMs) have become generally more capable and accurate across many tasks, some fundamental sources of unreliability remain in their behavior. One key limitation is their inconsistency at reporting the the same information when prompts are changed. In this paper, we consider the discrepancy between a model’s generated answer and their own verification of that answer, the generator-validator gap. We define this gap in a more stringent way than prior work: we expect correlation of scores from a generator and a validator over the entire set of candidate answers. We show that according to this measure, a large gap exists in various settings, including question answering, lexical semantics tasks, and next-word prediction. We then propose RankAlign, a ranking-based training method, and show that it significantly closes the gap by $31.8\%$ on average, surpassing all baseline methods. Moreover, this approach generalizes well to out-of-domain tasks and lexical items. [^1] Our code is available at https://github.com/juand-r/rankalign.
@@ 1 Introduction
However, this issue is not easily resolved in the face of epistemic uncertainty. Suppose that a model places probability mass over several answers to a question. Which of these answers should the model validate as correct? For a model to be consistent, the validator should not be more confident than the generator: in fact, the generator and validator’s confidences should be correlated.
In this paper, we introduce a new formulation of the generator-validator gap (G-V gap) that considers scores of the entire set of candidate answers at a time. We argue that achieving correlation between generator and validator scores across all answers, across all examples, is what is necessary for models to be consistent. Ideally, we want the model to refrain from generating responses that they disagree with discriminatively; we also expect the model to consistently assess answers even when they are less likely to generate the answers, a setting which arises when LMs are used as judges to evaluate arbitrary responses (Shi et al., 2024; Wang et al., 2024a; Li et al., 2024a). We show empirically that these correlations are low for existing open-source LLMs across a range of problems, including question answering, probing for lexical semantic knowledge (hypernymy and synonymy), and next-word prediction.
@@ 2 Problem Formulation
Giving the answer to a question and identifying whether a proposed answer is correct are two conceptually (Campbell, 1960) and computationally (Garey & Johnson, 1979) distinct problems. In the first, generative (generator) case, the LM must select an answer from among a combinatorially large set of options. In the second, discriminative (validator) case, the solution (or set of possible solutions) is presented along with the question, and the LM must select from among a small set of options, such as indicating if an answer is correct or incorrect. LMs (Hu & Frank, 2024; Li et al., 2024b) can give conflicting answers under generator and validator versions of the same question. This discrepancy is known as the Generator-Validator gap (Li et al., 2024b) or task demand gap between production and forced choice (Hu & Frank, 2024).
Figure 1 shows an example of how this gap can arise. When asked whether poodles are mammals, Gemma-2-2B prefers Yes over No, whereas mammal is a less likely continuation of ‘A poodle is a kind of’ than is the case for other examples with a strong preference for Yes, ranking 554th out of all examples. The opposite error can also occur. For example, fruit is an extremely likely continuation for ‘Complete the sentence: Olives are a kind of’ (in fact, it is the most likely token), even though the LM generates No when prompted with ‘Do you think olives are fruit?’
@@
Figure 2: Generator and validator log-odds for Gemma-2-2B for hypernym prediction (Pearson $\rho=0.764).$
@@
We measure the consistency between generator and validator across the range of answer choices $\{{y_{A_{i}}}\}_{i}$ via the Pearson correlation ($\rho$-all) of their log-odds $\{l_{G}(z,y_{A_{i}})\}_{z,i}$ and $\{l_{V}(z,y_{A_{i}})\}_{z,i}$. Since this correlation will be higher when generators and validators are both accurate (e.g., as in Figure 2) , we also evaluate the Pearson correlation between $l_{G}$ and $l_{V}$ restricted to only the set of positive examples $\mathcal{P}$ ($\rho$-pos), or negative examples $\mathcal{N}$ ($\rho$-neg). [^3] In some datasets (e.g., most QA datasets) only positive labels are available since all answers in the dataset are correct, in which case $\mathcal{N}=\emptyset$.
@@ 3 Training to Improve G-V Correlation
The surface form of the ranking loss resembles DPO (Rafailov et al., 2023) but has fundamental differences. It pushes the LMs to establish preferences between fixed completions given pairs of prompts, rather than pairs of completions given fixed prompts. DPO also involves comparison with the reference model to prevent the LM from straying too far from the initialization (Rafailov et al., 2024). We perform an ablation by adding $p_{ref}$ to $\mathcal{L}_{G2V}$ and $\mathcal{L}_{V2G}$ in Appendix §C.2 to explore the effect of adding a reference model in our setting.
@@
[…]
@@ 4.4 Models
We evaluate all methods on three models: Gemma-2-2B, Llama-3.2-3B and Llama-3.2-3B-Instruct. We test Llama-3.2-3B-Instruct on TriviaQA and lexical substitution, which are harder for non-instruction-tuned models than Hypernymy or LAMBADA. We detail the hyperparameters and other training setting in Appendix §A.
@@ 5 Results
Results on the Gemma-2-2B model experiments are shown in Tables 1 and 2. Full results for the other models follow broadly similar trends and are given in Appendix §C. Our main objective is to improve the generator-validator correlation, but it is also important that generator and validator performance does not degrade substantially.
@@ 5.1 In-domain Experiments
Task Method $\rho$-pos R@0 A@100 MRR-P
LAMBADA Base 6.1 90.8 99.3 79.0
SFT 9.8 100 99.7 83.5
Consistency FT 17.1 100 99.7 83.1
DPO V2G -41.8 100 82.8 54.6
Ranking G2V 60.0 67.7 94.8 52.9
Ranking V2G 11.3 95.5 98.5 68.8
TriviaQA Base 19.4 63.7 88.6 52.8
SFT 18.4 99.9 90.5 59.3
Consistency FT 20.1 99.9 90.9 59.3
DPO V2G 18.2 100 85.7 50.3
Ranking G2V 56.8 39.8 44.5 9.6
Ranking V2G 29.9 99.9 71.6 17.0
RankAlign is effective at closing the G-V gap
On all models and tasks, RankAlign substantially enhances the correlations between generator and validator, with an average gain of $31.8$ on $\rho_{\text{all}}$. It outperforms all baselines both overall and within individual classes. For Hypernymy, RankAlign nearly closes the G-V gap, with a correlation $\rho$-all of 94.2 and per-class correlations of 87.5 and 89.0 for the positive and negative classes, respectively. On SWORDS (Table 1), RankAlign increases $\rho$-all from a base 58.4 to 76.6, and it roughly doubles $\rho$-pos and $\rho$-neg (from 32.8 and 36.5 to 67.6 and 60.4, respectively). Similar improvements from RankAlign also hold for Llama-3.2-3B
(Table 7).
@@
RankAlign-V2G closes the gap less well than RankAlign
Across all tasks and models (Tables 110), we see that RankAlign significantly outperforms RankAlign-V2G on improving correlations. It appears that there is not enough information in the ranking of validator probabilities to successfully shift the distribution for the generator. We hypothesize that the generator’s distribution may be more precisely calibrated. Base LLMs are often calibrated (Kadavath et al., 2022a), and the generator is essentially using the model as a base LLM. By contrast, the capacity to act as a discriminator is more heavily induced by alignment, and aligned LLMs may exhibit lower calibration (Zhu et al., 2023).
@@
5.2 Cross-domain Experiments
Hypernymy Gemma-2 17.8 32.7 43.9 13.4 19.1 14.4 -10.9 19.7
Llama-3.2 16.5 33.2 46.1 11.8 14.4 10.6 0.5 11.9
SWORDS Gemma-2 4.5 18.3 12.8 18.2 34.8 23.9 7.8 17.6
Llama-3.2 6.5 9.2 21.6 26.2 40.8 31.7 0.6 14.6
LAMBADA Gemma-2 -6.8 -8.7 -12.3 -26.5 -15.6 -20.9 57.8 -21.4
Llama-3.2 -0.6 -2.8 -0.2 -5.7 -5.2 -3.6 45.2 -15.5
TriviaQA Gemma-2 -20.3 -12.7 -19.4 -18.6 -2.2 -18.3 1.9 37.4
Llama-3.2 -0.3 3.9 7.5 15.3 22.4 13.3 -0.6 50.7
Table 4: Cross-dataset evaluation showing the difference in $\rho$ scores (between the RankAlign trained model and the base LM) when training on a dataset (row) and evaluating on target datasets (column). Scores greater than 0 indicate that RankAlign generalizes across tasks. Values in gray are the in-domain results derived from Tables 1, 2, 7, 8.
@@ 5.2 Cross-domain Experiments
Generalization across tasks
We next evaluate the correlations of RankAlign in a cross-domain fashion, i.e., training them on one dataset and evaluating them on another. While correlations are naturally lower than training and evaluating on the same dataset, we consider our method to have generalized if the correlations exceed those of the original base model. The deltas in correlations between the fine-tuned models and the base models across domains are given in Table 4.  [^6] Full results comparing how well methods generalize from SWORDS to Hypernymy are shown in Table 13. We note that RankAlign also has a higher correlation than the baselines in this cross-domain setting, with generator and validator accuracy similar or higher than the base model.
@@ 6 Related Work
It is common to attribute propositional attitudes (Nelson, 2024) such as belief to LMs (Jiang et al., 2020; Kadavath et al., 2022b, i.a.). Several studies have discussed whether LMs can have beliefs and what normative constraints (Lin, 2024) would need to be enforced for the LM’s beliefs to be consistent (Hofweber et al., 2024; Hase et al., 2024; Fierro et al., 2024). The ranking perspective in this paper is also conceptually similar to Spohn’s ranking theory of belief (Spohn, 2009) which assigns ordinal ranks to propositions, where $A\prec B$ when credence in $B$ is higher than $A$.
@@ 7 Conclusion
In this paper, we present a new view of the generator-validator gap based on correlation between log-odds assigned under a generator and a validator. We describe a new method for training models to exhibit stronger correlation via a ranking loss between pairs of examples. Results show that our RankAlign significantly outperforms all baselines with mild task accuracy degradation, is robust to prompt variations, and generalizes well to unseen data and tasks. We believe future work can improve the generalization of these gains further and also seek to mechanistically understand the origins of these gaps, leading to new methods for more reliable language modeling.
@@ Acknowledgments
This work was supported by NSF CAREER Award IIS-2145280, the NSF AI Institute for Foundations of Machine Learning (IFML), a grant from Open Philanthropy, and Good Systems, [^8] https://goodsystems.utexas.edu/ a UT Austin Grand Challenge to develop responsible AI technologies.
@@ Appendix A Hyperparameters and other experimental details
Llama-3.2 Base 0 0 0 0

### v2

@@ Abstract
Although large language models (LLMs) have become more capable and accurate across many tasks, some fundamental sources of unreliability remain in their behavior. One key limitation is their inconsistency at reporting the same information when prompts are changed. In this paper, we consider the discrepancy between a model’s generated answer and their own verification of that answer, the generator-validator gap. We define this gap in a more stringent way than prior work: we expect correlation of scores from a generator and a validator over the entire set of candidate answers, i.e., candidate completions that could possibly arise during ordinary language use without breaking Gricean norms. We show that according to this measure, a large gap exists in various settings, including question answering, lexical semantics tasks, and next-word prediction. We then propose RankAlign, a ranking-based training method, and show that it significantly closes the gap, surpassing all baseline methods. Moreover, this approach generalizes well to out-of-domain tasks and lexical items. [^1] Our code is available at https://github.com/juand-r/rankalign.
@@ 1 Introduction
Suppose that a model places probability mass over several typical answers to a question. Which of these answers should the model validate as correct? For a model to be consistent, the generator’s probabilities should reflect the validator’s judgments and vice versa: the generator and validator’s confidences should be correlated.
In this paper, we introduce a new formulation of the generator-validator gap (G-V gap) that considers scores of the entire set of candidate answers at a time. Ideally, we want the model to refrain from generating responses that they disagree with discriminatively. We also expect the model to consistently assess answers even when they are less likely to generate the answers, a setting which arises when LMs are used as judges to evaluate arbitrary responses (Shi et al., 2024; Wang et al., 2024a; Li et al., 2024a) . Both desiderata will be made precise through the correlation of generator and validator scores, which can be thought of as reflecting model credences (degrees of belief). [^2] We take the intentional stance (Dennett, 1989) and do not address the issue of whether LMs actually have beliefs. It would be odd to have low credence for a proposition that one is unlikely to agree with when posed as a Yes/No question; some nuances to this issue will be discussed in §2.
We show empirically that these correlations are low for existing open-source LLMs across a range of problems, including question answering, probing for lexical semantic knowledge (hypernymy and synonymy), and next-word prediction.
@@ 2 Problem Formulation
Giving the answer to a question and identifying whether a proposed answer is correct are two conceptually (Campbell, 1960) and computationally (Garey & Johnson, 1979) distinct problems. In the first, generative (generator) case, the LM must select an answer from among a combinatorially large set of options. In the second, discriminative (validator) case, the solution (or set of possible solutions) is presented along with the question, and the LM must select from among a small set of options, such as indicating if an answer is correct or incorrect. LMs can give conflicting answers under generator and validator versions of the same question. This discrepancy is known as the Generator-Validator gap (Li et al., 2024b) or task demand gap between production and forced choice (Hu & Frank, 2024).
Figure 1 shows an example of how this gap can arise. When asked whether purses are clothing, Gemma-2-2B prefers Yes over No, whereas clothing is a less likely continuation of ‘Purses are a kind of’ than is the case for other examples with a strong preference for Yes, ranking 32rd out of all examples. The opposite error can also occur. For example, fruit is an extremely likely continuation for ‘Complete the sentence: Olives are a kind of’ (in fact, it is the most likely token), even though the LM generates No when prompted with ‘Do you think olives are fruit?’
@@
Desired behavior
It is clear that cases with high generator scores and low validator scores are unacceptable—the LM is contradicting itself (Li et al., 2024b). The reverse case is more subtle: it may be appropriate to accept propositions (high validator score) that are unlikely to be uttered (low generator score); e.g., “Dolphins are entities.”. This is due to the difference between believing and uttering a proposition: it is a violation of Gricean principles (Grice, 1975) to say things that are too trivial, specific, or obscure, regardless of one’s belief. For example, “Dolphins are entities“ is uninformative; “Dolphins are Odonoceti” is too specific. Such extreme Gricean violations are rare in our data due to the nature of our tasks (§4.1): concept pairs are derived from commonly-ocurring categories (hypernymy), human-generated synonyms in context, predictable cloze-style completions, or simple question answering. While the completions vary in terms of typicality (Murphy, 2004), they do not do so drastically. [^3] Moreover, experimental evidence (Hampton, 1998; 2007; Koriat & Sorka, 2015) shows that people’s judgment of membership and typicality are in fact correlated.
@@
Figure 2: Generator and validator log-odds for Gemma-2-2B for hypernym prediction (Pearson $\rho=0.76).$
Correlation of Log-Odds
@@
We measure the consistency between generator and validator across the range of natural (ecologically valid) answer choices $\{{y_{A_{i}}}\}_{i}$ via the Pearson correlation ($\rho$-all) of their log-odds $\{l_{G}(z,y_{A_{i}})\}_{z,i}$ and $\{l_{V}(z,y_{A_{i}})\}_{z,i}$. Since this correlation will be higher when generators and validators are both accurate (e.g., as in Figure 2) , we also evaluate the Pearson correlation between $l_{G}$ and $l_{V}$ restricted to only the set of positive examples $\mathcal{P}$ ($\rho$-pos), or negative examples $\mathcal{N}$ ($\rho$-neg). [^5] In some datasets (e.g., most QA datasets) only positive labels are available since all answers in the dataset are correct, in which case $\mathcal{N}=\emptyset$.
@@ 3 Training to Improve G-V Correlation
RankAlign-$\alpha$
Finally, we explore the convex combination of RankAlign and RankAlign-V2G losses:
$\mathcal{L}_{V2G+V2G}(p_{\theta})$ $=\alpha\mathcal{L}_{G2V}(p_{\theta})+(1-\alpha)\mathcal{L}_{V2G}(p_{\theta})$ (5)
in order to test whether aligning in both directions simultaneously can both reduce the G-V gap [^6] Future work could consider a loss that selectively chooses the alignment direction per sample, based on whether generator or validator are more accurate. and achieve high generator and validator accuracy. We set $\alpha=0.5$ in our experiments.
@@
The surface form of the ranking loss in RankAlign resembles DPO (Rafailov et al., 2023) but has fundamental differences. It pushes the LMs to establish preferences between fixed completions given pairs of prompts, rather than pairs of completions given fixed prompts. DPO also involves comparison with the reference model to prevent the LM from straying too far from the initialization (Rafailov et al., 2024). We perform an ablation by adding $p_{ref}$ to $\mathcal{L}_{G2V}$ and $\mathcal{L}_{V2G}$ in Appendix §C.2 to explore the effect of adding a reference model in our setting.
@@
To train with our RankAlign objective, we need to sample pairs of prompts $x_{G}$ and answers $A$ such that there is a clear ranking between them. We sample pairs with a minimum margin separation of $\delta$, i.e., such that ${\log(p_{LM}({A_{w}}\mid x_{G_{w}})})-{\log(p_{LM}({A_{l}}\mid x_{G_{l}})})\geq\delta$ for some $\delta>0$. [^7] Preliminary experiments with Gemma-2-2B across all tasks showed a drop in both correlations and accuracy when setting $\delta=0$. Pairs are sampled uniformly at random over the dataset and filtered based on this criterion. We also report an ablation where pairs are only sampled over the set of positive examples $\mathcal{P}$ in section §5. Other hyperparameters for training are described in Appendix §A.
@@ 4.4 Models
We evaluate all methods on four models: Gemma-2-2B, Llama-3.2-3B, Llama-3.2-3B-Instruct, and Llama-3.1-8B. We test Llama-3.2-3B-Instruct on TriviaQA and lexical substitution, which are harder for non-instruction-tuned models than Hypernymy or LAMBADA. We detail the hyperparameters and other training setting in Appendix §A.
@@ 5 Results
Results on the Gemma-2-2B model experiments are shown in Tables 2 and 1. Full results for the other models follow broadly similar trends and are given in Appendix §C. Our main objective is to improve the generator-validator correlation, but it is also important that generator and validator performance does not degrade substantially. In particular, when aligning the validator to the generator, the generator should be minimally impacted, and vice versa.
@@ 5.1 In-domain Experiments
Task Method $\rho$-pos R@0 A@100 MRR-P
LAMBADA Base 6.1 90.8 99.3 79.0
SFT 9.8 100 99.7 83.5
Consistency FT 17.1 100 99.7 83.1
DPO V2G -41.8 100 82.8 54.6
Ranking G2V 60.0 67.7 94.8 52.9
Ranking V2G 11.3 95.5 98.5 68.8
RankAlign-$\alpha$ 57.2 2.7 97.3 74.0
TriviaQA Base 19.4 63.7 88.6 52.8
SFT 18.4 99.9 90.5 59.3
Consistency FT 20.1 99.9 90.9 59.3
DPO V2G 18.2 100 85.7 50.3
Ranking G2V 56.8 39.8 44.5 9.6
Ranking V2G 29.9 99.9 71.6 17.0
RankAlign-$\alpha$ 73.1 75.2 80.0 36.9
RankAlign is effective at closing the G-V gap
On all models and tasks, RankAlign substantially enhances the correlations between generator and validator, with an average gain of $31.8$ on $\rho_{\text{all}}$. It outperforms all baselines both overall and within individual classes. For Hypernymy, RankAlign nearly closes the G-V gap, with a correlation $\rho$-all of 94.2 and per-class correlations of 87.5 and 89.0 for the positive and negative classes, respectively. On SWORDS (Table 2), RankAlign increases $\rho$-all from a base 58.4 to 76.6, and it roughly doubles $\rho$-pos and $\rho$-neg (from 32.8 and 36.5 to 67.6 and 60.4, respectively). Similar improvements from RankAlign also hold for Llama-3.2-3B and Llama-3.1-8B (Tables 7, 11).
@@
RankAlign variants
Across all tasks and models (Tables 210), we see that RankAlign significantly outperforms RankAlign-V2G on improving correlations, while RankAlign-$\alpha$ performs similarly. It appears that there is not enough information in the ranking of validator probabilities to successfully shift the distribution for the generator. We hypothesize that the generator’s distribution may be more precisely calibrated. Base LLMs are often calibrated (Kadavath et al., 2022a), and the generator is essentially using the model as a base LLM. By contrast, the capacity to act as a discriminator is more heavily induced by alignment, and aligned LLMs may exhibit lower calibration (Zhu et al., 2023).
5.2 Cross-domain Experiments
Generalization across tasks
We next evaluate the correlations of RankAlign in a cross-domain fashion, i.e., training them on one dataset and evaluating them on another. While correlations are naturally lower than training and evaluating on the same dataset, we consider our method to have generalized if the correlations exceed those of the original base model. The deltas in correlations between the fine-tuned models and the base models across domains are given in Table 3.  [^8] Full results comparing how well methods generalize from SWORDS to Hypernymy are shown in Table 15. We note that RankAlign also has a higher correlation than the baselines in this cross-domain setting, with generator and validator accuracy similar or higher than the base model. Additional results for RankAlign-$\alpha$ can be found in Table 16.
Hypernymy Gemma-2 17.8 32.7 43.9 13.4 19.1 14.4 -10.9 19.7
Llama-3.2 16.5 33.2 46.1 11.8 14.4 10.6 0.5 11.9
SWORDS Gemma-2 4.5 18.3 12.8 18.2 34.8 23.9 7.8 17.6
Llama-3.2 6.5 9.2 21.6 26.2 40.8 31.7 0.6 14.6
LAMBADA Gemma-2 -6.8 -8.7 -12.3 -26.5 -15.6 -20.9 57.8 -21.4
Llama-3.2 -0.6 -2.8 -0.2 -5.7 -5.2 -3.6 45.2 -15.5
TriviaQA Gemma-2 -20.3 -12.7 -19.4 -18.6 -2.2 -18.3 1.9 37.4
Llama-3.2 -0.3 3.9 7.5 15.3 22.4 13.3 -0.6 50.7
Table 3: Cross-dataset evaluation showing the difference in $\rho$ scores (between the RankAlign trained model and the base LM) when training on a dataset (row) and evaluating on target datasets (column). Scores greater than 0 indicate that RankAlign generalizes across tasks. Values in gray are the in-domain results derived from Tables 2, 1, 7, 8.
@@ 5.2 Cross-domain Experiments
Generalization across prompt formats
@@ 6 Related Work
It is common to attribute propositional attitudes (Nelson, 2024) such as belief to LMs (Jiang et al., 2020; Kadavath et al., 2022b, i.a.). Several studies have discussed whether LMs can have beliefs and what normative constraints (Lin, 2024) would need to be enforced for the LM’s beliefs to be consistent (Hofweber et al., 2024; Hase et al., 2024; Fierro et al., 2024). The ranking perspective in this paper is also conceptually similar to Spohn’s ranking theory of belief (Spohn, 2009) which assigns ordinal ranks to propositions. Finally, Gekhman et al. (2025) find that LLMs encode significantly more factual knowledge in their parameters than they express in their outputs.
@@ 7 Conclusion
In this paper, we present a new view of the generator-validator gap based on correlation between log-odds assigned under a generator and a validator. We describe a new method for training models to exhibit stronger correlation via a ranking loss between pairs of examples. Results show that our RankAlign significantly outperforms baselines with mild task accuracy degradation, is robust to prompt variations, and generalizes well to unseen data and tasks.
We believe future work can improve the generalization of these gains further and also seek to mechanistically understand the origins of these gaps, leading to new methods for more reliable language modeling. Other directions include extending the investigation to longer-form completions, which are found in LLM evaluations of summarization or long-form QA (Xu et al., 2023). On the theoretical side, it would be interesting to formulate the desired (ideal, rational) behavior relating generator and validator probabilities to a wider range of cases which violate Grice’s maxims. This forces one to grapple fully with the distinction between beliefs and speech acts. One approach to this could be to take inspiration from Rational Speech Act theory (Frank & Goodman, 2012) in order to model what a rational speaker should utter (rather than believe) under generator and validator settings.
@@ Acknowledgments
This work was supported by NSF CAREER Award IIS-2145280, the NSF AI Institute for Foundations of Machine Learning (IFML), a grant from Open Philanthropy, and Good Systems, [^11] https://goodsystems.utexas.edu/ a UT Austin Grand Challenge to develop responsible AI technologies. We thank the anonymous reviewers whose helpful feedback helped improve the work.
@@ Appendix A Hyperparameters and other experimental details
RankAlign-$\alpha$ 3491 3919 3203 3644
Llama-3.2-3B Base 0 0 0 0
@@
RankAlign-$\alpha$ - 3312 3573 -
Llama-3.1-8B Base 0 0 0 0
SFT 4497 1052 6000 8306
DPO G2V 3000 696 3000 4153
DPO V2G 2149 329 - -
Consistency FT 2364 510 2242 3214
RankAlign 3274 3116 3242 3484
RankAlign-V2G 4168 4505 3530 4598
RankAlign-$\alpha$ 3274 3166 3242 3484